Can you help me solve a system of equations by graphing (just explain the steps) for the following system. Can you give the x- and y-intercepts and the slope of each equation? Thanks.
x-3y=-14; 3x+8y=26

Answer by stanbon(75887)

x-3y=-14
x-intercept: Let y=0, then x=-14
y=intercept: Let x=0, then y=14/3
slope: solve for y=(x-14)/3
y= (1/3)x-(14/3)
So, slope is 1/3
3x+8y=26
x-intercept: Let y=0, then x= 26/3
y-intercept: Let x=0, then y=26/8
slope: solve for y=(-3x+26)/8
y= (-3/8)+(26/8)
So, slope is -3/8
Cheers,
stan H.